# Edinburg, Texas, USA

### v1

@@ 2.4 Effects of input correlations
Correlation or synchrony in neuronal activity can be described for any readout of brain activity. Here, we are concerned with the spiking activity of neurons.
@@
In the simplest way, correlation/synchrony refers to the coincident spiking of neurons, i.e., when two neurons spike together, they are firing in synchrony or are correlated. Neurons can be synchronous in their instantaneous activity, i.e., they spike together with some probability. However, it is also possible that a neuron’s spiking at time $t$ is correlated with the spikes of another neuron with a delay (time-delayed synchrony).
Origin of synchronous neuronal activity:
@@
  • Common inputs, i.e., two neurons are receiving input from the same sources. The degree of correlation of the shared inputs is proportional to their output correlation.
@@
  • Pooling from the same sources. Neurons do not share the same input neurons but are receiving inputs from neurons which themselves are correlated.
  • Neurons are connected to each other (uni- or bi-directionally): This will only give rise to time-delayed synchrony. Neurons could also be connected via gap-junctions.
  • Neurons have similar parameters and initial conditions.
When neurons spike together, they can have a stronger impact on downstream neurons. Synapses in the brain are sensitive to the temporal correlations (i.e., delay) between pre- and postsynaptic activity, and this, in turn, can lead to the formation of functional neuronal networks - the basis of unsupervised learning.
Synchrony implies a reduction in the dimensionality of the system. In addition, correlations, in many cases, can impair the decoding of neuronal activity.
A simple model to study the emergence of correlations is to inject common inputs into a pair of neurons and measure the output correlation as a function of the fraction of common inputs.
@@
Here, we are going to investigate the transfer of correlations by computing the correlation coefficient of spike trains recorded from two unconnected HH neurons, which received correlated inputs.
The input current to HH neuron $i=1,2$ is
$I_{i}=\mu_{i}+\sigma_{i}(\sqrt{1-c}\xi_{i}+\sqrt{c}\xi_{c}),$ (32)
@@
where $\mu_{i}$ is the temporal average of the current. The Gaussian white noise $\xi_{i}$ is independent for each neuron, while $\xi_{c}$ is common to all neurons. The variable $c$ $(0\leq c\leq 1)$ controls the fraction of common and independent inputs. $\sigma_{i}$ shows the variance of the total input.
The sample correlation coefficient between two input currents, $I_{i}$ and $I_{j}$, is defined as the sample covariance of $I_{i}$ and $I_{j}$ divided by the square root of the sample variance of $I_{i}$ multiplied by the square root of the sample variance of $I_{j}$. We consider the following equations:
$r_{ij}$ $=\frac{\mathrm{cov}(I_{i},I_{j})}{\sqrt{\mathrm{var}(I_{i})}\sqrt{\mathrm{var}(I_{i})}},$ (33)
$\mathrm{cov}(I_{i},I_{j})$ $=\sum_{k=1}^{L}(I_{i}^{k}-\bar{I}_{i})(I_{j}^{k}-\bar{I}_{j}),$ (34)
$\mathrm{var}(I_{i})$ $=\sum_{k=1}^{L}(I_{i}^{k}-\bar{I}_{i})^{2},$ (35)
where $\bar{I}_{i}$ is the sample mean, $k$ is the time bin, and $L$ is the length of $I$. This means that $I_{i}^{k}$ is current at $i$ at time $k\cdot dt$. Note that the equations above are not accurate for sample covariances and variances as they should be additionally divided by $L-1$. We have dropped this term because it cancels out in the sample correlation coefficient formula.
@@ 2.5 STDP in neuromorphic systems and other applications
With many current and potential applications, STDP is often thought of as an unsupervised brain-like learning mechanism for spiking neural networks (SNNs) that, among other things, has attracted significant attention from the neuromorphic hardware community [41, 29]. Its ability to mimic biological learning processes makes STDP highly relevant for various applications, including pattern recognition and sensory processing, real-time pattern recognition, stabilized supervised STDP, and synchronization in neural networks. Many models have been proposed to investigate the role of STDP mechanisms across these applications. For example, the human brain is recognized as the most complex entity in the known universe [19]. At the microcircuit level, neuronal cells are organized into layers with various connectivity motifs. While the information processing mechanisms at this level remain not fully understood, investigating these motifs—particularly about STDP—is essential for gaining insights into biological learning mechanisms and the emergence of intelligence. STDP is an unsupervised brain-like learning rule implemented in many SNNs and neuromorphic chips. However, a significant performance gap exists between ideal model simulation and neuromorphic implementation. STDP is implemented in SNNs and neuromorphic chips, serving as an unsupervised learning rule crucial for mimicking brain-like information processing [18]. Its applications include noisy spatiotemporal spike pattern detection, which is particularly effective even with low-resolution synaptic efficacy in neuromorphic implementations. This capability enhances performance in various computational tasks, making STDP relevant for advancing neuromorphic hardware [13, 14]. STDP is used to train an efficient Spiking Auto-Encoder that leverages asynchronous sparse spikes for input reconstruction, denoising, and classification, achieving superior performance with significantly fewer spikes compared to state-of-the-art methods. On the other hand, STDP has significant applications in bio-plausible meta-learning models, particularly in enhancing the adaptability and efficiency of machine-learning systems. By incorporating STDP and Reward-Modulated STDP, these models reflect biological learning mechanisms and enable quick learning in low-data scenarios. This approach is particularly useful for preventing catastrophic forgetting in meta-learning tasks, allowing the model to retain previously acquired knowledge while learning new tasks. Additionally, STDP facilitates the application of these models in spike-based neuromorphic devices, improving their performance in tasks such as few-shot classification and advancing AI systems’ capabilities to mimic human-like learning [23]. STDP is employed to train a Spiking Auto-Encoder that efficiently performs input reconstruction, denoising, and classification with minimal spike usage, showcasing enhanced performance on image datasets while maintaining competitiveness against other learning approaches [46]. STDP is also applied through memristors as synapses to enable in situ learning and inference in SNNs. This approach addresses the computational challenges associated with implementing STDP in hardware, allowing for efficient weight modulation that enhances speed and reduces power consumption. The integration of STDP facilitates real-time pattern recognition by employing a winner-takes-all mechanism within the SNN architecture. The proposed design significantly improves performance metrics, including power, energy, and accuracy, enabling the classification of 50 million images per second [22]. STDP is employed in the Stabilized Supervised STDP learning rule to enhance the classification layer of SNNs, integrating unsupervised STDP for feature extraction and improving performance on image recognition tasks [16]. Furthermore, STDP is utilized to investigate the configurations needed to achieve robust synchronization in neural networks, with findings that could inform the design of neuromorphic circuits for improved information processing and transmission through synchronization phenomena [51]. Overall, despite the challenges in scaling STDP for deeper networks and larger tasks, its biological relevance and versatility highlight its significance in advancing artificial intelligence and neural computation. STDP holds the potential to enhance the performance of robotic systems by allowing them to learn from their environments in real time, reinforcing its critical role in the development of intelligent systems.
@@ 3 Numerical results
In particular, we use the coupled SDEs-ODEs system (2.8)–(2.20) that describes the dynamics of the STN membrane potential. As we have mentioned in the previous section, we will focus on the effects of Gaussian white noise input current together with the random refractory periods on the STN cell membrane potential.
@@
In order to switch from healthy conditions to Parkinsonian conditions in the basal ganglia model, we consider a decrease in the current $I_{\text{app}}$ applied to the STN. In particular, we have $I_{\text{app}}=33$ (pA) for a healthy STN cell and $I_{\text{app}}=23$ (pA) for a Parkinsonian STN cell (see, e.g., [37]). Therefore, an STN cell in the case of injected current input $I_{\text{app}}=33$ (pA) results in a healthy STN cell, while an STN cell in the case of injected current $I_{\text{app}}=23$ (pA) is considered as a PD-affected STN cell.
@@
Figure 2: [Color online] Membrane potential profile of direct input current and direct refractory period of healthy cells.
@@
In particular, we look at the Fig. 2, where we have plotted the time evolution of the STN cell membrane potential with direct input current and direct refractory period of healthy cells. The neuron spikes are different between the neuron spikes in the case of healthy cells and with PD cells presented in Figs. 2-3. There are missing spikes in the time interval of $[0,35]$ in Fig. 3.
Figure 3: [Color online] Membrane potential profile of direct input current and direct refractory period of PD.
Figure 4: [Color online] Membrane potential profile of direct input current and direct refractory period of healthy cells. Parameters: $\sigma=1,\sigma_{\text{ref}}=1$.
@@
In Fig. 4, in the case of healthy cell, we added the random input current and random refractory period to the system. We observe that there are fluctuations in the time evolution of the membrane potential, the total excitatory synaptic conductance as well as the trace of the number of postsynaptic spikes over the timescale $\tau_{-}$. In the case of PD-affected cells with random input current and random refractory period in Fig. 5, we see that the time evolution of the cell membrane potential is more unstable. There are also fluctuations in the trace of the number of postsynaptic conductance and the total excitatory synaptic conductance.
Figure 5: [Color online] Membrane potential profile of random input current and random refractory period of PD. Parameters: $\sigma=1,\sigma_{\text{ref}}=1$.
Figure 6: [Color online] Membrane potential profile of random input current and random refractory period of PD cells with DBS.
@@
In Fig. 6, we consider a DBS input to the system, there are still fluctuations in the time evolution of the membrane potential and the trace of the number of postsynaptic spikes over the timescale $\tau_{-}$. In what follows, we also look at the corresponding ISI distributions and the spike irregularity profiles of the cases presented in Fig. 7.
@@
Figure 7: Spike irregularity profiles in the case with the additive noise input current and the ISI distribution. From the top to the bottom, first 3 rows: ISI distributions with direct refractory period. Second 3 rows: ISI distributions with random refractory period. Last 3 rows: ISI distributions with random refractory period and DBS. Parameters: with $\sigma=1,\sigma_{\text{ref}}=1$ in the left panel, and $\sigma=10,\sigma_{\text{ref}}=5$ in the right panel.
In Fig. 7, we present spiking irregularity profiles in the case of an additive noise input current, along with the ISI distribution for healthy and Parkinson’s disease (PD)-affected cells. The variability in ISI is typically measured by its coefficient of variation, $CV_{\text{ISI}}$. Our analysis focuses on values at $I_{\text{app}}=23$ pA and $I_{\text{app}}=33$ pA. We first determine the average current injection values, $I_{\text{average}}=[22,34]$ pA. The ISI is computed by identifying the spike times and calculating the differences between them. The $CV_{\text{ISI}}$ is then defined as per equation (31).
@@
In Fig. 8, the plot of input correlation versus output correlation is referred to as the correlation transfer function of neurons. The results indicate that output correlation is lower than input correlation and output correlation varies linearly with input correlation.
When input correlations do not affect the neuron’s capacity, output correlation remains independent of both the mean and standard deviation. However, we observe that increasing the standard deviation of the random refractory period in both healthy and PD cells leads to a linear increase in output correlations (see, for example, [9, 28]).

### v2

@@ 2.4 Effects of input correlations
Correlation or synchrony in neuronal activity refers to the relationship between the firing patterns of different neurons, and it can be measured in various ways. In this context, we focus on the spiking activity of neurons. At its simplest, correlation or synchrony refers to the coincident spiking of neurons-that is, when two neurons spike together, they are said to be firing in synchrony or to be correlated [52]. Neurons can be synchronous in their instantaneous activity, meaning they spike together with some probability. However, synchrony can also occur with a time delay, where the spiking of one neuron at time $t$ is correlated with the spiking of another neuron at a later time (time-delayed synchrony). Origins of synchronous neuronal activity consist of:
@@
  • Common inputs: Neurons receiving input from the same sources tend to have correlated activity. The degree of correlation in their inputs influences the degree of correlation in their outputs.
@@
  • Pooling from correlated sources: Neurons may not share the same input neurons but could receive inputs from other neurons that are themselves correlated.
  • Direct connections: Neurons connected to each other (either unidirectionally or bidirectionally) can exhibit time-delayed synchrony. Gap-junctions between neurons can also facilitate synchrony.
  • Similar properties: Neurons with similar intrinsic parameters and initial conditions may also exhibit synchronous behavior.
When neurons fire together, their coordinated activity can have a stronger influence on downstream neurons. Synapses are sensitive to the temporal correlations between presynaptic and postsynaptic spikes, and this sensitivity plays a crucial role in shaping functional neuronal networks?essential for unsupervised learning. While synchrony can reduce the system’s dimensionality, strong correlations may sometimes impair the decoding of neuronal activity. A simple model to study the emergence of correlations involves injecting common inputs into two neurons and measuring the resulting output correlation as a function of the fraction of shared inputs.
@@
In this study, we will investigate how correlations are transferred by calculating the correlation coefficient of spike trains recorded from two unconnected HH neurons that received correlated inputs [52, 53]. The input current to HH neurons $i=1,2$ is given by:
$I_{i}=\mu_{i}+\sigma_{i}(\sqrt{1-c}\,\xi_{i}+\sqrt{c}\,\xi_{c}),$
@@
where $\mu_{i}$ represents the temporal average of the current. The Gaussian white noise $\xi_{i}$ is independent for each neuron, while $\xi_{c}$ is common to all neurons. The parameter $c$ (where $0\leq c\leq 1$) controls the proportion of common and independent inputs, and $\sigma_{i}$ is the variance of the total input. The sample correlation coefficient between the input currents $I_{i}$ and $I_{j}$ is defined as the sample covariance of $I_{i}$ and $I_{j}$ divided by the product of the square roots of their sample variances. Specifically, we use the following equations:
$r_{ij}=\frac{\mathrm{cov}(I_{i},I_{j})}{\sqrt{\mathrm{var}(I_{i})}\,\sqrt{\mathrm{var}(I_{j})}},$
$\mathrm{cov}(I_{i},I_{j})=\sum_{k=1}^{L}(I_{i}^{k}-\bar{I}_{i})(I_{j}^{k}-\bar{I}_{j}),$
$\mathrm{var}(I_{i})=\sum_{k=1}^{L}(I_{i}^{k}-\bar{I}_{i})^{2},$
where $\bar{I}_{i}$ is the sample mean of $I_{i}$, $k$ is the time bin index, and $L$ is the total number of samples. Here, $I_{i}^{k}$ represents the current at neuron $i$ at time $k\cdot dt$. Note that these formulas for covariance and variance are not fully accurate as they should be divided by $L-1$ for sample estimates. However, we omit this correction because it cancels out in the calculation of the sample correlation coefficient.
@@ 2.5 STDP in neuromorphic systems and other applications
With many current and potential applications, STDP is often thought of as an unsupervised brain-like learning mechanism for spiking neural networks (SNNs) that, among other things, has attracted significant attention from the neuromorphic hardware community [54, 55]. Its ability to mimic biological learning processes makes STDP highly relevant for various applications, including pattern recognition and sensory processing, real-time pattern recognition, stabilized supervised STDP, and synchronization in neural networks. Many models have been proposed to investigate the role of STDP mechanisms across these applications. For example, the human brain is recognized as the most complex entity in the known universe [56]. At the microcircuit level, neuronal cells are organized into layers with various connectivity motifs. While the information processing mechanisms at this level remain not fully understood, investigating these motifs?particularly about STDP?is essential for gaining insights into biological learning mechanisms and the emergence of intelligence. STDP provides plasticity rules that depend on spikes. As such, they are unsupervised learning rules commonly used in spiking SNNs and neuromorphic chips to emulate brain-like information processing. However, a significant performance gap remains between ideal model simulations and their neuromorphic implementations. STDP is implemented in SNNs and neuromorphic chips, serving as an unsupervised learning rule crucial for mimicking brain-like information processing [57]. Its applications include noisy spatiotemporal spike pattern detection, which is particularly effective even with low-resolution synaptic efficacy in neuromorphic implementations. This capability enhances performance in various computational tasks, making STDP relevant for advancing neuromorphic hardware [58, 59]. STDP is used to train an efficient Spiking Auto-Encoder that leverages asynchronous sparse spikes for input reconstruction, denoising, and classification, achieving superior performance with significantly fewer spikes compared to state-of-the-art methods. On the other hand, STDP has significant applications in bio-plausible meta-learning models, particularly in enhancing the adaptability and efficiency of machine-learning systems. By incorporating STDP and Reward-Modulated STDP, these models reflect biological learning mechanisms and enable quick learning in low-data scenarios. This approach is particularly useful for preventing catastrophic forgetting in meta-learning tasks, allowing the model to retain previously acquired knowledge while learning new tasks. Additionally, STDP facilitates the application of these models in spike-based neuromorphic devices, improving their performance in tasks such as few-shot classification and advancing AI systems’ capabilities to mimic human-like learning [60]. STDP is employed to train a Spiking Auto-Encoder that efficiently performs input reconstruction, denoising, and classification with minimal spike usage, showcasing enhanced performance on image datasets while maintaining competitiveness against other learning approaches [61]. STDP is also applied through memristors as synapses to enable in situ learning and inference in SNNs. This approach addresses the computational challenges associated with implementing STDP in hardware, allowing for efficient weight modulation that enhances speed and reduces power consumption. The integration of STDP facilitates real-time pattern recognition by employing a winner-takes-all mechanism within the SNN architecture. The proposed design significantly improves performance metrics, including power, energy, and accuracy, enabling the classification of 50 million images per second [62]. STDP is employed in the Stabilized Supervised STDP learning rule to enhance the classification layer of SNNs, integrating unsupervised STDP for feature extraction and improving performance on image recognition tasks [63]. Furthermore, STDP is utilized to investigate the configurations needed to achieve robust synchronization in neural networks, with findings that could inform the design of neuromorphic circuits for improved information processing and transmission through synchronization phenomena [64].
STDP can be regarded as a key learning rule in biological neural networks, and its relevance has been explored in neuromorphic systems for a variety of applications. STDP enables synaptic modifications based on the relative timing of pre- and post-synaptic spikes, which is critical for encoding information and adapting to dynamic environments. Recent advancements in flexible neuromorphic electronics have made it possible to integrate artificial synapses and neurons that replicate the functionality of their biological counterparts, providing a platform for computing, soft robotics, and neuroprosthetics. These systems, which emulate synaptic behaviors and exhibit learning capabilities, are particularly promising for future applications in health monitoring and the Internet of Things [65]. Furthermore, neuromorphic computing algorithms, which draw inspiration from biological learning rules like STDP, are becoming central to the development of more efficient, adaptive, and scalable computing technologies. The ongoing research in neuromorphic hardware and algorithms is setting the stage for new opportunities in both machine learning and real-world applications such as cognitive computing and neuroprosthetics [66].
Overall, despite the challenges in scaling STDP for deeper networks and larger tasks, its biological relevance and versatility highlight its significance in advancing artificial intelligence and neural computation. STDP holds the potential to enhance the performance of robotic systems by allowing them to learn from their environments in real time, reinforcing its critical role in the development of intelligent systems.
@@ 3 Numerical results
In particular, we use the coupled SDEs-ODEs system (2.8)?(2.20) that describes the dynamics of the STN membrane potential. As we have mentioned in the previous section, we will focus on the effects of Gaussian white noise input current together with the random refractory periods on the STN cell membrane potential.
@@
In order to switch from healthy conditions to Parkinsonian conditions in the STN model, we consider a decrease in the current $I_{\text{app}}$ applied to the STN. In particular, we have $I_{\text{app}}=33$ (pA) for a healthy STN cell and $I_{\text{app}}=23$ (pA) for a Parkinsonian STN cell (see, e.g., [27]). Therefore, an STN cell in the case of injected current input $I_{\text{app}}=33$ (pA) results in a healthy STN cell, while an STN cell in the case of injected current $I_{\text{app}}=23$ (pA) is considered as a PD-affected STN cell.
@@
Figure 2: [Color online] Figures showing the evolution of synaptic conductance and membrane potential over time. (A): Membrane potential (V) as a function of time (in milliseconds) (2.3). (B): Presynaptic timing (M) dynamics (4). (C): Average excitatory synaptic conductance ($\bar{g}$) updates for multiple synapses (6). (D): Excitatory synaptic conductance ($g_{E}$) (8). (E): Example of presynaptic spike trains for excitatory neurons. (F): The effective conductance or strength of synaptic transmission for excitatory neurons ($P_{ex}$) (5). (G): Example of presynaptic spike trains for inhibitory neurons. (H): The effective conductance or strength of synaptic transmission for inhibitory neurons ($P_{in}$) (5). These panels illustrate the time-dependent changes in membrane potential and synaptic activity profile of direct input current and direct refractory period of healthy cells.
@@
In particular, we look at the Fig. 2 A, where we have plotted the time evolution of the STN cell membrane potential with direct input current and direct refractory period of healthy cells. The neuron spikes are different between the neuron spikes in the case of healthy cells and with PD cells presented in Figs. 2-3. There are missing spikes in the time interval of $[0,35]$ in Fig. 3 A.
Figure 3: [Color online] Figures showing the evolution of synaptic conductance and membrane potential over time. (A): Membrane potential (V) as a function of time (in milliseconds) (2.3). (B): Postsynaptic timing (M) dynamics (4). (C): Average excitatory synaptic conductance ($\bar{g}$) updates for multiple synapses (6). (D): Excitatory synaptic conductance ($g_{E}$) (8). (E): Example of presynaptic spike trains for excitatory neurons. (F): The effective conductance or strength of synaptic transmission for excitatory neurons ($P_{ex}$) (5). (G): Example of presynaptic spike trains for inhibitory neurons. (H): The effective conductance or strength of synaptic transmission for inhibitory neurons ($P_{in}$) (5). These panels illustrate the time-dependent changes in membrane potential and synaptic activity profile of direct input current and direct refractory period of PD.
Figure 4: [Color online] Figures showing the evolution of synaptic conductance and membrane potential over time. (A): Membrane potential (V) as a function of time (in milliseconds) (2.3). (B): Postsynaptic timing (M) dynamics (4). (C): Average excitatory synaptic conductance ($\bar{g}$) updates for multiple synapses (6). (D): Excitatory synaptic conductance ($g_{E}$) (8). (E): Example of presynaptic spike trains for excitatory neurons. (F): The effective conductance or strength of synaptic transmission for excitatory neurons ($P_{ex}$) (5). (G): Example of presynaptic spike trains for inhibitory neurons. (H): The effective conductance or strength of synaptic transmission for inhibitory neurons ($P_{in}$) (5). These panels illustrate the time-dependent changes in membrane potential and synaptic activity profile of direct input current and direct refractory period of healthy cells. Parameters: $\sigma=1,\sigma_{\text{ref}}=1$.
@@
In Fig. 4, in the case of healthy cell, we added the random input current and random refractory period to the system. We observe that there are fluctuations in the time evolution of the membrane potential (Fig. 4 A), the total excitatory synaptic conductance (Fig. 4 C) as well as the trace of the number of postsynaptic spikes over the timescale $\tau_{-}$ (Fig. 4 B). In the case of PD-affected cells with random input current and random refractory period in Fig. 5, we see that the time evolution of the cell membrane potential is more unstable (Fig. 5 A). There are also fluctuations in the trace of the number of postsynaptic conductance (Fig. 5 C) and the total excitatory synaptic conductance (Fig. 5 B).
Figure 5: [Color online] Figures showing the evolution of synaptic conductance and membrane potential over time. (A): Membrane potential (V) as a function of time (in seconds). (B): Presynaptic timing (M) dynamics. (C): Average excitatory synaptic conductance ($\bar{g}$) updates for multiple synapses (6). (D): Excitatory synaptic conductance ($g_{E}$) (8). (E): Example of presynaptic spike trains for excitatory neurons. (F): Probability of presynaptic spike transmission for excitatory neurons ($P_{ex}$). (G): Example of presynaptic spike trains for inhibitory neurons. (H): Probability of presynaptic spike transmission for inhibitory neurons ($P_{in}$). These panels illustrate the time-dependent changes in membrane potential and synaptic activity profile of random input current and random refractory period of PD. Parameters: $\sigma=1,\sigma_{\text{ref}}=1$.
Figure 6: [Color online] Figures showing the evolution of synaptic conductance and membrane potential over time. (A): Membrane potential (V) as a function of time (in seconds). (B): Presynaptic timing (M) dynamics. (C): Average excitatory synaptic conductance ($\bar{g}$) updates for multiple synapses (6). (D): Excitatory synaptic conductance ($g_{E}$) (8). (E): Example of presynaptic spike trains for excitatory neurons. (F): Probability of presynaptic spike transmission for excitatory neurons ($P_{ex}$). (G): Example of presynaptic spike trains for inhibitory neurons. (H): Probability of presynaptic spike transmission for inhibitory neurons ($P_{in}$). These panels illustrate the time-dependent changes in membrane potential and synaptic activity profile of random input current and random refractory period of PD cells with DBS.
@@
In Fig. 6, we consider a DBS input to the system, there are still fluctuations in the time evolution of the membrane potential (Fig. 6 A) and the trace of the number of postsynaptic spikes over the timescale $\tau_{-}$ (Fig. 6 B). In what follows, we also look at the corresponding ISI distributions and the spike irregularity profiles of the cases presented in Fig. 7.
@@
Figure 7: Spike irregularity profiles in the case with the additive noise input current and the ISI distribution. From the top to the bottom, (A) and (B): ISI distributions with direct refractory period. (C) and (D): ISI distributions with random refractory period. (E) and (F): ISI distributions with random refractory period and DBS. Parameters: with $\sigma=1,\sigma_{\text{ref}}=1$ in (A), (C) and (E); and $\sigma=10,\sigma_{\text{ref}}=5$ in (B), (D) and (F).
In Fig. 7, we present spiking irregularity profiles in the case of an additive noise input current, along with the ISI distribution for healthy and Parkinson?s disease (PD)-affected cells. The variability in ISI is typically measured by its coefficient of variation, $CV_{\text{ISI}}$. Our analysis focuses on values at $I_{\text{app}}=23$ pA and $I_{\text{app}}=33$ pA. We first determine the average current injection values, $I_{\text{average}}=[22,34]$ pA. The ISI is computed by identifying the spike times and calculating the differences between them. The $CV_{\text{ISI}}$ is then defined as per equation (29).
@@
The results presented in Fig. 8 are obtained by simulating the network response for varying levels of input correlation (see e.g. subsection 2.4). For each input correlation, we compute the output cross-correlation across multiple trials, using a set of predefined model parameters and spike trains for the presynaptic neurons. The plot shows how the input correlation influences the output correlation, reflecting the impact of input correlation on network activity. In Fig. 8, the plot of input correlation versus output correlation is referred to as the correlation transfer function of neurons. The results indicate that output correlation is lower than input correlation and output correlation varies linearly with input correlation.
When input correlations do not affect the neuron?s capacity, output correlation remains independent of both the mean and standard deviation of the refractory period. However, we observe that increasing the standard deviation of the random refractory period in both healthy and PD cells leads to an increase in output correlations, particularly at lower input correlation values (below 0.2). This increase is linear in nature, meaning that as the input correlation increases within this range, the output correlation also increases in a proportional manner. Specifically, this linear trend is most pronounced when input correlations are small, and it reflects how variability in the refractory period amplifies the output correlation as input correlation values approach lower thresholds (see, for example, [67, 68]).